4. Suppose that P{X = i, Y = j} = c(i+j) for nonnegative integers i and j with i+j < 3; otherwise, the probability is zero. (I need help on parts e-f
Answer #1

i,j Cam take o, 1, 2, 3 :. 0,0) = (ico, j=0) аluе. 2 3 | P13) | с 16 по 1 о с І. Яс 2 | e 3 3 | 3 o Рx x ) с 6c ч яр с о o o☺ Elx] = Ex pcx) X = 0+66 +1466 +2456+3+ 30 = 66 + 10 C +90 325€ = 2 +1.25 ETY] - Ey Pey) DeretnSo2+29olto900 = = [scill=352